# Efficiently Computing Equilibria in Budget-Aggregation Games

### v2

@@ 5 Leontief utilities
Brandt et al. (2023, 2025) introduced Leontief preferences for contexts where alternatives share characteristics of complements — meaning that agents only benefit when they receive all of them (in a specific ratio), rather than being able to substitute one for another. For each agent $i$, let $A^{+}_{i}:=\{x\in A\colon v_{i,x}>0\}$ denote the set of alternatives the agent values positively. Leontief preferences are represented by
@@
Definition 5.1 (Brandt et al. (2025)).
Given a distribution $\delta\in\Delta$, agent $i$’s set of critical alternatives is defined as
$T_{\delta,i}\coloneqq\arg\min_{x\in A^{+}_{i}}\frac{\delta(x)}{v_{i,x}}\text{.}$
@@
Lemma 5.2 (Brandt et al. (2023, 2025), Lemma 1).
@@
If the agents’ valuations $v_{i,x}$ are natural numbers, then the equilibrium distribution $\delta^{*}$ and its utility profile $u^{*}$ are rational-valued.
Moreover, their binary encoding length is bounded by a polynomial function of the binary encoding length of $v_{i,x}$ and $n$.
@@
For each $i\in N$, let $T_{i}$ be a non-empty set of alternatives. Consider the following linear program (LP), with variables $u_{i}$ (for $i\in N$), $d_{x}$ (for $x\in A$), and $d_{i,x}$ (for $i\in N$ and $x\in A$):
@@
We can even give an explicit bound on the representation length of the equilibrium distribution $\delta^{*}$. For that, we construct an undirected graph where vertices correspond to alternatives, and there is an edge between $x,y\in A$ if and only if there exists an agent $i\in N$ with $x,y\in T_{\delta^{*},i}$. Each component of that graph can be considered separately, as the sum of contributions to that component equals the fraction of agents for whom parts of the component are critical alternatives.
@@
This gives a system of linear equations where each $\delta^{*}_{x_{j}}$ can be written in terms of $\delta^{*}_{x_{1}}$. Specifically, $\delta^{*}_{x_{2}}$ only requires the two valuations $v_{i_{1},x_{1}}$ and $v_{i_{1},x_{2}}$, $\delta^{*}_{x_{3}}$ requires at most four valuations ($v_{i_{1},x_{1}},v_{i_{1},x_{2}}$ and $v_{i_{2},x_{2}},v_{i_{2},x_{3}}$ ) , and so on. Considering this “worst case” in terms of representation length together with $\sum_{x\in A^{\prime}}\delta^{*}_{x}=\lvert N^{\prime}\rvert/n$, $\delta^{*}_{x_{1}}$ can be written as a fraction with numerator $\lvert N^{\prime}\rvert/n$ and denominator $1+v_{i_{1},x_{2}}/v_{i_{1},x_{1}}+(v_{i_{1},x_{2}}/v_{i_{1},x_{1}})\cdot(v_{i_{2},x_{3}}/v_{i_{2},x_{2}})+\dots$. This results in a binary encoding length of $\log_{2}((\lvert N^{\prime}\rvert/n)\cdot{v_{\max}}^{m-1})$ for the numerator and $\log_{2}(m\cdot{v_{\max}}^{m-1}))$ for the denominator where we upper-bounded $m^{\prime}$ by $m$, and $v_{i,x}$ by $v_{\max}$.
@@
It remains to handle the case that all linear constraints are satisfied, whereas the nonlinear constraint is violated. That is, we have
$\prod_{i=1}^{n}(u^{\prime}_{i})<{z_{0}}^{n}.$ (1)
@@
Lemma 5.6 allows us to apply Lemma 5.5 to $S(z_{0})$. We can now compute the equilibrium distribution by applying binary search to $z_{0}$, in the following way.
@@
Moreover, Theorem 5.3 is applicable to other utility models where NEs coincide with the ones for Leontief utilities, e.g., Cobb-Douglas utility functions (Brandt et al., 2025).
@@ 6 Binary symmetric separable utilities
where $g_{i,x}\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ are continuous and weakly increasing functions.
@@
For example, setting $g_{i}(\delta_{x})=\delta_{x}$ for all $i\in N$ corresponds to linear preferences with binary valuations $v_{i,x}\in\{0,1\}$ (see, e.g., Bogomolnaia et al., 2005). Choosing $g_{i}(\delta_{x})=\ln(\delta_{x})$ leads to Cobb-Douglas utilities with binary weights. However, note that our model allows the $g_{i}$’s to differ among agents.
The following example shows that for such general preferences, NEs do not always exist.
@@
Intuitively, Agent $1$ wants the entire budget to be distributed as equally as possible, whereas Agent $2$ prefers to focus on one alternative.
Therefore, we treat separately the two cases where all $g_{i}$’s are strictly concave (Section 6.1) and strictly convex (Section 6.2).
6.1 Strictly concave case
In this section, all $g_{i}$’s are assumed to be strictly concave, i.e., $g_{i}(\lambda\delta+(1-\lambda)\delta^{\prime})>\lambda g_{i}(\delta)+(1-\lambda)g_{i}(\delta^{\prime})$ for all different $\delta,\delta^{\prime}\in\Delta$ and $\lambda\in(0,1)$.
Theorem 6.2 (Brandt et al. (2025), Proposition 9).
@@ 6.1 Strictly concave case
For binary valuations, Brandt et al. (2025) stated two different possibilities to compute an overall equilibrium distribution via linear programming. This distribution can be decomposed into an NE using the same linear program described at the end of Section 5.
@@
6.2 Strictly convex case
In this section, all $g_{i}$’s are assumed to be strictly convex, i.e., $\lambda g_{i}(\delta)+(1-\lambda)g_{i}(\delta^{\prime})>g_{i}(\lambda\delta+(1-\lambda)\delta^{\prime})$ for all different $\delta,\delta^{\prime}\in\Delta$ and $\lambda\in(0,1)$.
Note that then each utility function $u_{i}(\delta)$, seen as an $m$-dimensional real-valued function, is convex, as it is a sum of strictly convex functions. Furthermore, it is strictly convex if the domain is restricted to the $\lvert A_{i}\rvert-$dimensional subspace spanned by agent $i$’s approved alternatives.
Let $M_{\delta,i}\coloneqq\arg\max_{x\in A_{i}}\delta_{x}$ be the set of alternatives with maximal share under the distribution $\delta$, which are approved by agent $i$.
@@ 6.2 Strictly convex case
The distribution $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if, every agent $i$’s distribution $\delta^{*}_{i}$ is a best response and therefore a solution of the optimization problem
@@
The objective function is strictly convex on this domain and the feasible region is the ($\lvert A_{i}\rvert-1$)-dimensional simplex spanned by those strategies where agent $i$ places her entire budget on a single approved alternative. So every maximizer is a vertex of this simplex, which proves that $\lvert\operatorname{supp}(\delta_{i})\rvert=1$. By adding $1/n$ to an alternative $x\in A_{i}$ in $\delta^{*}-\delta^{*}_{i}$, agent $i$’s utility increases by $g(\delta^{*}_{x}-\delta^{*}_{i,x}+1/n)-g(\delta^{*}_{x}-\delta^{*}_{i,x})$. This increase is maximal if and only if $x$ is an alternative with maximal share under the distribution $\delta^{*}-\delta^{*}_{i}$, i.e., $x\in M_{\delta^{*}-\delta^{*}_{i},i}$ since $g$ is strictly convex. This shows that $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if every agent plays a pure strategy, i.e., $\lvert\operatorname{supp}(\delta^{*}_{i})\rvert=1$, and $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{\delta^{*}-\delta^{*}_{i},i}$ for all $i\in N$. Thus, $\lvert M_{\delta^{*},i}\rvert=1$ and $\operatorname{supp}(\delta^{*}_{i})=M_{\delta^{*},i}$ for all $i\in N$.
@@
$\delta_{y}-\delta_{i,x}=\delta_{y}\leq\delta_{x}-1/n=\delta_{x}-\delta_{i,x}$
@@
We use the following greedy approach to construct an NE. Let $N^{\prime}\subseteq N$ denote the set of agents who have not yet allocated their budgets. In each step, select an alternative $x^{*}\in\arg\max_{x\in A}\sum_{i\in N^{\prime}:x\in A_{i}}1/n$, i.e., $x^{*}$ maximizes the aggregated share of the remaining agents who approve it. Each of these agents then allocates her entire budget to $x^{*}$, and is excluded from $N^{\prime}$. The procedure is repeated until no agents remain.
@@
In fact, the set of NEs is even discrete, as there are at most finitely many NE.
@@
7 $\ell_{1}$ preferences
A widely studied utility model in the context of budget-aggregation is based on $\ell_{1}$ preferences, also referred to as $\ell_{1}$ disutilities (Lindner, 2011; Goel et al., 2019; Freeman et al., 2019). The existing literature mostly focuses on the axiomatic investigation of mechanisms (Brandt et al., 2024; Elkind et al., 2024), in particular moving phantom mechanisms (Freeman et al., 2021; de Berg et al., 2024).
For $\ell_{1}$ preferences, each agent is assumed to have a peak allocation — a preferred distribution of the budget — and her utility is determined by the negative $\ell_{1}$ distance between any given allocation and her peak. Formally, let $p_{i}\in\Delta(A)$ denote the peak of agent $i$, then her utility for a given allocation $\delta\in\Delta(A)$ is defined as
@@ 7 $\ell_{1}$ preferences
These utilities capture the behavior of agents who derive value from receiving resources up to a certain satiation point, beyond which additional resources provide no further benefit. Specifically, the utility is linear in the allocation up to the agent’s peak, and constant thereafter — reflecting satiation once one’s desired threshold is met.
We say an alternative $x$ is oversupplied for agent $i$ if $\delta_{x}>p_{i,x}$. Additionally, let $M_{i}^{>}(\delta)$ be the set of alternatives that are oversupplied for agent $i$ given an allocation $\delta$. Similarly, we say $x$ is undersupplied for $i$ if $\delta_{x}<p_{i,x}$ and $M_{i}^{<}(\delta)$ is the set of undersupplied alternatives. We show that a strategy profile constitutes an NE when no agent contributes to one of her oversupplied alternatives.
@@
For $\ell_{1}$ preferences, a strategy profile $(\delta_{i})_{i\in N}$ is a Nash equilibrium if and only if $supp(\delta_{i})\subseteq M_{i}^{\leq}(\delta)$ for all $i\in N$.
@@
$\Rightarrow$: Assume that there is some agent $i$ with $\delta_{i,x}>0$ for some alternative $x$ which is oversupplied, i.e. $\delta_{x}>p_{i,x}$. Redistributing $\delta_{i,x}$ to some undersupplied alternative, say $x^{\prime}$, strictly increases the utility of agent $i$, showing that the strategy profile $(\delta_{i})_{i\in N}$ is not an NE.
$\Leftarrow$: Suppose $(\delta_{i})_{i\in N}$ satisfies $supp(\delta_{i})\subseteq M_{i}^{\leq}(\delta)$ for all $i\in N$. Since the marginal utility decrease in $x$ equals the marginal utility increase in $x^{\prime}$ for all $x,x^{\prime}\in M_{i}^{\leq}(\delta)$, agent $i$ cannot strictly increase her utility by redistributing her budget $\delta_{i}$. Therefore, $u_{i}(\delta)=\max_{\delta_{i}^{\prime}\in S_{i}}u_{i}(\delta-\delta_{i}+\delta_{i}^{\prime})$, establishing that $(\delta_{i})_{i\in N}$ is an NE. ∎
The existence of an NE for $\ell_{1}$ disutilities follows from the more general existence result by Debreu (1952) for quasi-concave utilities. Given our above characterization, it is unsurprising that the NE is not unique in the case of $\ell_{1}$ preferences since the characterization solely depends on the support of the agents’ individual distributions.
@@
With the help of the above example, we can show the following claim.
@@
We present a polynomial-time algorithm. Let each agent $i$ have a budget, denoted by $b_{i}$. Initially, this budget is set to $b_{i}=1/n$ for all $i\in N$. Iterate over all alternatives $x\in A$ in an arbitrary order. For a given alternative $x$, order all agents in a descending order of their peaks for this alternative, i.e., $p_{i_{1},x}\geq p_{i_{2},x}\geq\dots\geq p_{i_{n},x}$. When it is agent $i$’s turn, she allocates $\min(b_{i},p_{i,x}-\delta_{x})$ on alternative $x$. Afterwards, $b_{i}$ is decreased and $\delta_{x}$ is increased by this amount. Once agent $i$ has spent her entire share ($b_{i}=0$), she is deleted from the set of considered agents. The ordering of the agents ensures that $x$ does not become oversupplied to an agent who already contributed to $x$ later on, i.e., no agent contributes to an oversupplied alternative. When no agent wants to contribute to alternative $x$, we go to the next alternative $x^{\prime}$ and reorder the remaining agents according to their peaks ($p_{i_{1},x^{\prime}}\geq p_{i_{2},x^{\prime}}\dots\geq p_{i_{n},x^{\prime}}$). Then, we let them again sequentially contribute to this alternative.
@@
This construction provides an efficient algorithm to obtain an NE given a set of optimal distributions of the agents. Additionally, we can also show whether a given overall distribution can be represented as individual distributions that constitute an NE.

### v3

@@ 5 Leontief utilities
Brandt et al. (2023, 2025a) introduced Leontief preferences for contexts where alternatives share characteristics of complements — meaning that agents only benefit when they receive all of them (in a specific ratio), rather than being able to substitute one for another. For each agent $i$, let $A^{+}_{i}:=\{x\in A\colon v_{i,x}>0\}$ denote the set of alternatives the agent values positively. Leontief preferences are represented by
@@
Definition 5.1 (Brandt et al. (2025a)).
Given a distribution $\delta\in\Delta(A)$, agent $i$’s set of critical alternatives is defined as
$T_{\delta,i}\coloneqq\arg\min_{x\in A^{+}_{i}}\frac{\delta_{x}}{v_{i,x}}\text{.}$
@@
Lemma 5.2 (Brandt et al. (2023, 2025a), Lemma 1).
@@
If the agents’ valuations $v_{i,x}$ are natural numbers, then the equilibrium distribution $\delta^{*}$ and its utility profile $u^{*}$ are rational-valued.   Moreover, their binary encoding length is bounded by a polynomial function of the binary encoding length of $v_{i,x}$ and $n$.
@@
For each $i\in N$, let $T_{i}$ be a non-empty set of alternatives. Consider the following linear program (LP), with variables $u_{i}$ (for $i\in N$), $d_{x}$ (for $x\in A$), and $d_{i,x}$ (for $i\in N$ and $x\in T_{i}$):
@@
We can give an explicit bound on the representation length of the equilibrium distribution $\delta^{*}$. For that, we construct an undirected graph where vertices correspond to alternatives, and there is an edge between $x,y\in A$ if and only if there exists an agent $i\in N$ with $x,y\in T_{\delta^{*},i}$. Each component of that graph can be considered separately, as the sum of contributions to that component equals the fraction of agents for whom parts of the component are critical alternatives.
@@
This gives a system of linear equations where each $\delta^{*}_{x_{j}}$ can be written in terms of $\delta^{*}_{x_{1}}$. Specifically, $\delta^{*}_{x_{2}}$ only requires the two valuations $v_{i_{1},x_{1}}$ and $v_{i_{1},x_{2}}$, $\delta^{*}_{x_{3}}$ requires at most four valuations ($v_{i_{1},x_{1}},v_{i_{1},x_{2}}$ and $v_{i_{2},x_{2}},v_{i_{2},x_{3}}$ ) , and so on. Considering this “worst case” in terms of representation length together with $\sum_{x\in A^{\prime}}\delta^{*}_{x}=\lvert N^{\prime}\rvert/n$, $\delta^{*}_{x_{1}}$ can be written as a fraction with numerator $\lvert N^{\prime}\rvert/n$ and denominator $1+v_{i_{1},x_{2}}/v_{i_{1},x_{1}}+(v_{i_{1},x_{2}}/v_{i_{1},x_{1}})\cdot(v_{i_{2},x_{3}}/v_{i_{2},x_{2}})+\dots$. This results in a binary encoding length of $\log_{2}((\lvert N^{\prime}\rvert/n)\cdot{v_{\max}}^{m-1})$ for the numerator and $\log_{2}(m\cdot{v_{\max}}^{m-1})$ for the denominator, where we used $m^{\prime}\leq m$, and $v_{i,x}\leq v_{\max}$.
@@
It remains to handle the case that all linear constraints are satisfied, whereas the nonlinear constraint is violated, i.e.,
$\prod_{i=1}^{n}u^{\prime}_{i}<{z_{0}}^{n}.$ (1)
@@
Lemma 5.6 allows us to apply Lemma 5.5 to $S(z_{0})$. We can now compute the equilibrium distribution by applying binary search to $z_{0}$, as follows.
@@
Moreover, Theorem 5.3 is applicable to other utility models where NEs coincide with the ones for Leontief utilities, e.g., Cobb-Douglas utility functions (Brandt et al., 2025b).
@@ 6 Binary symmetric separable utilities
where $g_{i,x}\colon\mathbb{R}_{\geq 0}\to\mathbb{R}$ are continuous and weakly increasing.
@@
For example, setting $g_{i}(\delta_{x})=\delta_{x}$ for all $i\in N$ corresponds to linear preferences with binary valuations $v_{i,x}\in\{0,1\}$ (see, e.g., Bogomolnaia et al., 2005). Choosing $g_{i}(\delta_{x})=\ln(\delta_{x})$ leads to Cobb-Douglas utilities with binary weights. However, note that our model allows the $g_{i}$’s to differ among agents.   The following example shows that for such general preferences, NEs do not always exist.
@@
Intuitively, Agent $1$ wants the entire budget to be distributed as equally as possible, whereas Agent $2$ prefers to focus on one alternative.   Therefore, we treat separately the two cases where all $g_{i}$’s are strictly concave (Section 6.1) and strictly convex (Section 6.2).
6.1 Strictly concave utilities
In this section, all $g_{i}$’s are assumed to be strictly concave, i.e., $g_{i}(\lambda\delta+(1-\lambda)\delta^{\prime})>\lambda g_{i}(\delta)+(1-\lambda)g_{i}(\delta^{\prime})$ for all different $\delta,\delta^{\prime}\in\Delta(A)$ and $\lambda\in(0,1)$.
Theorem 6.2 (Brandt et al. (2025a), Proposition 9).
@@ 6.1 Strictly concave utilities
For binary valuations, Brandt et al. (2025a) stated two different possibilities to compute an overall equilibrium distribution via linear programming. This distribution can be decomposed into an NE using the same linear program described at the end of Theorem 5.3.
@@
6.2 Strictly convex utilities
In this section, all $g_{i}$’s are assumed to be strictly convex, i.e., $\lambda g_{i}(\delta)+(1-\lambda)g_{i}(\delta^{\prime})>g_{i}(\lambda\delta+(1-\lambda)\delta^{\prime})$ for all different $\delta,\delta^{\prime}\in\Delta(A)$ and $\lambda\in(0,1)$.
Note that in this case, each utility function $u_{i}(\delta)$, seen as an $m$-dimensional real-valued function, is convex, as it is a sum of strictly convex functions. Furthermore, it is strictly convex if the domain is restricted to the $\lvert A_{i}\rvert-$dimensional subspace spanned by agent $i$’s approved alternatives.
Let $M_{\delta,i}\coloneqq\arg\max_{x\in A_{i}}\delta_{x}$ be the set of alternatives with maximum share under the distribution $\delta$ that are approved by agent $i$.
@@ 6.2 Strictly convex utilities
The distribution $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if every agent $i$’s distribution $\delta^{*}_{i}$ is a best response and therefore a solution of the optimization problem
@@
The objective function is strictly convex on this domain, and the feasible region is the ($\lvert A_{i}\rvert-1$)-dimensional simplex spanned by those strategies where agent $i$ places her entire budget on a single approved alternative. So every maximizer is a vertex of this simplex, which proves that $\lvert\operatorname{supp}(\delta_{i})\rvert=1$. By adding $1/n$ to an alternative $x\in A_{i}$ in $\delta^{*}-\delta^{*}_{i}$, agent $i$’s utility increases by $g_{i}(\delta^{*}_{x}-\delta^{*}_{i,x}+1/n)-g_{i}(\delta^{*}_{x}-\delta^{*}_{i,x})$. This increase is maximal if and only if $x$ is an alternative with maximal share under the distribution $\delta^{*}-\delta^{*}_{i}$, i.e., $x\in M_{\delta^{*}-\delta^{*}_{i},i}$ since $g_{i}$ is strictly convex. This proves that $(\delta^{*}_{i})_{i\in N}$ is an NE if and only if every agent plays a pure strategy, i.e., $\lvert\operatorname{supp}(\delta^{*}_{i})\rvert=1$, and $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{\delta^{*}-\delta^{*}_{i},i}$ for all $i\in N$. Thus, $\lvert M_{\delta^{*},i}\rvert=1$ and $\operatorname{supp}(\delta^{*}_{i})=M_{\delta^{*},i}$ for all $i\in N$.
@@
$\delta_{y}-\delta_{i,y}=\delta_{y}\leq\delta_{x}-1/n=\delta_{x}-\delta_{i,x}$
@@
We use the following greedy approach to construct an NE $(\delta_{i}^{*})_{i\in N}$.  [^8] This algorithm is equivalent to an algorithm called “maximum payment rule” (Aziz et al., 2025), which was recently presented as a rule for approximately-fair budget distribution for agents with binary linear utilities. Let $N^{\prime}\subseteq N$ denote the set of agents who have not yet allocated their budgets. In each step, select an alternative $x^{*}\in\arg\max_{x\in A}\sum_{i\in N^{\prime}:x\in A_{i}}1/n$, i.e., $x^{*}$ maximizes the aggregated share of the remaining agents who approve it. Each of these agents then allocates her entire budget to $x^{*}$, and is excluded from $N^{\prime}$. The procedure is repeated until no agents remain.
@@
In fact, the number of NE is finite and can be bounded explicitly.
@@
7 $\ell_{1}$ disutilities
A widely studied utility model in the context of budget-aggregation is based on $\ell_{1}$ preferences, also referred to as $\ell_{1}$ disutilities (Lindner, 2011; Goel et al., 2019; Freeman et al., 2019). The existing literature mostly focuses on the axiomatic investigation of mechanisms (Brandt et al., 2024; Elkind et al., 2026), in particular, moving phantom mechanisms (Freeman et al., 2021; de Berg et al., 2024).
For $\ell_{1}$ preferences, each agent is assumed to have a peak allocation — a preferred distribution of the budget — and her utility is determined by the negative $\ell_{1}$ distance between any given allocation and her peak. Formally, let $p_{i}=(p_{i,x})_{x\in A}\in\Delta(A)$ denote the peak of agent $i$, then her utility for a given allocation $\delta\in\Delta(A)$ is defined as
@@ 7 $\ell_{1}$ disutilities
These utilities capture the behavior of agents who derive value from receiving resources up to a certain satiation point, beyond which additional resources provide no further benefit. Specifically, the utility is linear in the allocation up to the agent’s peak, and constant thereafter — reflecting satiation once the desired threshold is met.
We say an alternative $x$ is oversupplied for agent $i$ if $\delta_{x}>p_{i,x}$. Additionally, let $M_{i}^{>}(\delta)$ be the set of alternatives that are oversupplied for agent $i$ given an allocation $\delta$. Similarly, an alternative $x$ is undersupplied for agent $i$ if $\delta_{x}<p_{i,x}$ and $M_{i}^{<}(\delta)$ is the set of undersupplied alternatives. Define $M_{i}^{\leq}(\delta^{*})$ accordingly. We show that a strategy profile constitutes an NE when no agent contributes to one of her oversupplied alternatives.
@@
For $\ell_{1}$ preferences, a strategy profile $(\delta^{*}_{i})_{i\in N}$ is a Nash equilibrium if and only if $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{i}^{\leq}(\delta^{*})$ for all $i\in N$.
@@
$\Rightarrow$: Assume that there is some agent $i$ with $\delta^{*}_{i,x}>0$ for some alternative $x$ which is oversupplied, i.e. $\delta^{*}_{x}>p_{i,x}$. Redistributing $\delta^{*}_{i,x}$ to some undersupplied alternative, say $x^{\prime}$, strictly increases the utility of agent $i$, showing that the strategy profile $(\delta^{*}_{i})_{i\in N}$ is not an NE.
$\Leftarrow$: Suppose $(\delta^{*}_{i})_{i\in N}$ satisfies $\operatorname{supp}(\delta^{*}_{i})\subseteq M_{i}^{\leq}(\delta^{*})$ for all $i\in N$. Since the marginal utility decrease in $x$ equals the marginal utility increase in $x^{\prime}$ for all $x,x^{\prime}\in M_{i}^{\leq}(\delta^{*})$, agent $i$ cannot strictly increase her utility by redistributing her budget $\delta_{i}$. Therefore, $u_{i}(\delta^{*})=\max_{\delta_{i}^{\prime}\in S_{i}}u_{i}(\delta^{*}-\delta^{*}_{i}+\delta_{i}^{\prime})$, establishing that $(\delta^{*}_{i})_{i\in N}$ is an NE. ∎
The existence of an NE for $\ell_{1}$ disutilities follows from the more general existence result by Debreu (1952) for quasi-concave utilities. Given our above characterization, it is not surprising that NEs do not yield a unique overall distribution in the case of $\ell_{1}$ preferences, since the characterization depends solely on the support of the agents’ individual distributions.
@@
With the above example, we can show the following.
@@
We present a polynomial-time algorithm. Let each agent $i$ have a budget, denoted by $b_{i}$. Initially, this budget is set to $b_{i}=1/n$ for all $i\in N$. Iterate over all alternatives $x\in A$ in an arbitrary order. For a given alternative $x$, order all agents in a descending order of their peaks for this alternative, i.e., $p_{i_{1},x}\geq p_{i_{2},x}\geq\dots\geq p_{i_{n},x}$. When it is agent $i$’s turn, she allocates $\min(b_{i},\max(0,p_{i,x}-\delta_{x}))$ on alternative $x$. Afterwards, $b_{i}$ is decreased and $\delta_{x}$ is increased by this amount. Once agent $i$ has spent her entire share ($b_{i}=0$), she is deleted from the set of considered agents. The ordering of the agents ensures that $x$ does not become oversupplied to an agent who already contributed to $x$ later on, i.e., no agent contributes to an oversupplied alternative. When no agent wants to contribute to alternative $x$, we go to the next alternative $x^{\prime}$ and reorder the remaining agents according to their peaks ($p_{i_{1},x^{\prime}}\geq p_{i_{2},x^{\prime}}\dots\geq p_{i_{n},x^{\prime}}$). Then, we let them again sequentially contribute to this alternative.
@@
This construction provides an efficient algorithm for obtaining an NE given a set of optimal distributions of the agents. Additionally, we can also show whether a given overall distribution can be represented as individual distributions that constitute an NE.